OpenStudy (anonymous):

What is the quotient: (2x^3 – 15x^2 – 10x + 16) ÷ (x – 8) ?

OpenStudy (anonymous):

(2x^3 – 15x^2 – 10x + 16) ÷ (x – 8) = 2x² +x -2 2x^3 - 16x^2 ------------ 1 x² - 10 x + 16 x² - 8x ------------ -2x + 16 -2x +16 ----- 0

OpenStudy (anonymous):

Test: (2x² + x - 2) * (x - 8) = 2x³ - 16x² + x² - 8x - 2x + 16 = 2x³ -15x² -10x +16

OpenStudy (anonymous):

thanks would u mind helping me along the lines of this test aswell?

OpenStudy (anonymous):

i think that's gonna be no problem for me :)

OpenStudy (anonymous):

ok thank you so much im working out this problem real quick its like the same

OpenStudy (anonymous):

alright :)

OpenStudy (anonymous):

What are the zeros of the polynomial function: f(x) = x^3 + x^2 – 20x ?

OpenStudy (anonymous):

ok i never undersstood zeros and pellet

OpenStudy (anonymous):

f(x) = x^3 + x^2 – 20x f(x) = x* ( x² + x - 20) x1 = 0 ( Product Zero sentence) f(x) = x² + x - 20 20 = x² + x 20 = x* ( x + 1 ) x = 20 (Product Zero Sentence) 0 = x + 1 x = -1 x1,2,3 = 0,0, -1

OpenStudy (anonymous):

thats not one of the answers D: